OpenStudy (abbles):

If f and f^-1 are both differentiable for all x, with f(3) = 5 and f'(3) = 7 , then write the equation of a line tangent to the graph of f^-1

OpenStudy (abbles):

I literally have no clue..

zepdrix (zepdrix):

f(3)=5, which means f^-1(5)=3, ya?

zepdrix (zepdrix):

\[\large\rm y=mx+b\]We're trying to create some linear function which is tangent to the inverse function at ... oh you didn't specify a location. I assume it's at x=3, yes? You can find the `slope` by using your formula for derivative inverse,\[\large\rm m=(f^{-1})'(3)=\frac{1}{f'(f^{-1}(3))}\]

zepdrix (zepdrix):

Oh maybe this tangent line is being created at x=5. That would make more sense. Fill in the details when you come back :D

OpenStudy (abbles):

Would the tan line be 1/7 for the slope?

OpenStudy (abbles):

Would the coordinate be (5, 3)

OpenStudy (abbles):

Would the equation be y - 3 = 1/7(x-5)

zepdrix (zepdrix):

Ooo yes good job!\[\large\rm (f^{-1})'(5)=\frac{1}{f'(f^{-1}(5))}=\frac{1}{f'(3)}=\frac{1}{7}=m\]And since \(\large\rm f^{-1}(5)=3\) we have the coordinate (5,3) for the inverse function.